### v1

@@ Abstract
On-board sensors of autonomous vehicles can be obstructed, occluded, or limited by restricted fields of view, complicating downstream driving decisions. Intelligent roadside infrastructure perception systems, installed at elevated vantage points, can provide wide, unobstructed intersection coverage, supplying a complementary information stream to autonomous vehicles via vehicle-to-everything (V2X) communication. However, conventional 3D object-detection algorithms struggle to generalize under the domain shift introduced by top-down perspectives and steep camera angles. We introduce a 2.5D object detection framework, tailored specifically for infrastructure roadside-mounted cameras. Unlike conventional 2D or 3D object detection, we employ a prediction approach to detect ground planes of vehicles as parallelograms in the image frame. The parallelogram preserves the planar position, size, and orientation of objects while omitting their height, which is unnecessary for most downstream applications. For training, a mix of real-world and synthetically generated scenes is leveraged. We evaluate generalizability on a held-out camera viewpoint and in adverse-weather scenarios absent from the training set. Our results show high detection accuracy, strong cross-viewpoint generalization, and robustness to diverse lighting and weather conditions. Model weights and inference code are provided at:
@@ I Introduction
Autonomous vehicles primarily rely on on-board sensors such as cameras and Lidar for environmental perception. However, these sensors are inherently limited by occlusions and narrow fields of view, posing challenges in dynamic urban environments where relevant objects may not be visible. This is particularly critical at traffic intersections, where autonomous driving systems must explicitly handle non-observable regions to ensure safe driving maneuvers [1]. Unlike vehicle-mounted sensor systems, roadside infrastructure-based methods leverage elevated vantage points to capture wider fields of view with reduced occlusions, sometimes utilizing V2X communication for information transmission [2, 3]. However, these elevated viewpoints also introduce new challenges. Although 3D object detection models perform well with on-board vehicle sensors, they often fail to generalize to elevated viewing angles. Furthermore, to reduce costs, intelligent infrastructure typically employs monocular cameras rather than Lidar, limiting the range of applicable 3D object detection methods. Compared to the abundance of datasets for autonomous vehicle perception [4, 5, 6], available datasets for infrastructure-based perception are limited, necessitating efficient data collection and training strategies. Current state-of-the-art methods exhibit several limitations, such as assuming uniform vehicle dimensions, enforcing alignment with lane directions and past trajectories, or relying on planar road surfaces [7, 8, 9, 10, 11, 12]. To address these challenges, we present a real-time–capable 2.5D object detection framework. It can be trained efficiently with minimal real-world data, generalizes across diverse camera angles and weather conditions, and outperforms previous approaches.
@@ II-A Camera-based Object Detection
Camera-based object detection is typically categorized into 2D and 3D detection, both aiming to distinguish relevant objects from the background. Standard 2D object detectors detect objects by predicting their $x,y$ position, width, and height in pixel coordinates to generate axis-aligned boxes, which are then classified. Most baseline architectures are trained and evaluated on the MS COCO dataset [14] with large architectures, oftentimes based on Detection Transformer (DETR) [15], leading benchmark performance [16, 17, 18]. A separate line of research prioritizes inference speed to enable real-time object detection without significantly sacrificing accuracy. Here, YOLO-based models [19, 20, 21, 22, 23, 24, 25, 26] and RT-DETR methods [27, 28, 29] are among the most commonly employed approaches.
3D camera-based object detection detects objects in real-world 3D coordinates. These methods typically predict a rectangular cuboid, defined by the object’s $x,y,z$ position, width, height, and length. The object’s orientation, expressed in Euler angles, is also predicted. Because most 3D object detection datasets in the traffic domain require additional Lidar data for ground truth annotations, they are typically constructed using data collected from vehicles such as KITTI [4], nuScenes [5], or Waymo [6]. In this domain, the majority of detected objects, vehicles and pedestrians, are situated on a mostly planar road surface, allowing for the common simplification of assuming zero pitch and roll angles. Monocular 3D object detection is an inherently ill-posed problem, as 2D images introduce scale ambiguities that must be approximated by inferring depth and learning geometric priors of frequently encountered objects [30].
One monocular approach, M3D-RPN [31], combines 2D and 3D region proposal networks using 3D anchors to create strong priors correlating 2D scale with 3D depth. For orientation, only the yaw angle is predicted, and it is further refined by re-projecting the 3D cuboid into the image plane and rotating objects to maximize overlap with 2D detections. Another approach, FCOS3D [13], extends the anchor-free FCOS [32] 2D object detection framework to monocular 3D object detection by projecting the 3D object centers onto the image plane and decoupling 2D attributes from depth and orientation estimates.
@@ II-B Object Detectors for Roadside Infrastructure
[7], [8], [9], and [10] rely on 2D object detections projected from image frame into world coordinates. They assume fixed object sizes and estimate orientation from past trajectories. These approaches tackle the ill-posed nature of monocular 3D detection by relying on geometric approximations and precise camera calibration. For example, Rezai et al. [10] employ approximately 70 images per camera, along with top-view satellite imagery, to perform inverse perspective mapping [33] and project 2D detections into 3D. They assign a fixed width and length to each object class, which is used for depth estimation of 2D object detections. To estimate the object’s heading (yaw angle), they track past detections and approximate the orientation based on the object’s motion trajectory. Since trajectory-based orientation estimation fails for static objects, they also assume parallel heading relative to known road lanes. They evaluate their method on 2D datasets [34, 35, 36] and the non-public Leeds [10] dataset.
@@
Zhu et al. [12], similarly to Rezai et al. [10], annotate keypoints in satellite and roadside camera images and compute a homography mapping images of roadside cameras into a bird’s-eye view (BEV) assuming planar roads. 2D object detection is subsequently applied to these BEV images, estimating rotated bounding boxes, which correspond to the yaw-angle orientation. However, transforming roadside images into BEV space creates large distortions.
@@ III-A Dataset Creation
For training the proposed 2.5D object detector, datasets containing annotated parallelograms are required. But, to the best of our knowledge, no publicly available dataset is dedicated to ground-plane estimation from roadside cameras. Popular datasets like MIO-TCD [34], UA-DETRAC [35], and GRAM-RTM [36] provide only 2D bounding box information and therefore lack the geometric information needed to recover ground planes. Conversely, 3D datasets are suitable in principle, as the required ground plane information can be calculated by projecting the four bottom vertices of each 3D bounding box into the image plane. Since the Rope3D [38] and Leeds [10] datasets are inaccessible, we examine the TUMTraf [3, 40] dataset for potential use. While camera viewpoints are suitable, we observe that a significant portion of annotations are inaccurate (cf. Figure 4), limiting the dataset’s direct applicability to training our model.
@@
To obtain reliable data, we recorded a new dataset with the intelligent roadside infrastructure of the Test Area Autonomous Driving Baden-Württemberg (TAF-BW) [7, 41, 42]. We captured 3,633 images using seven roadside cameras. Sampling images at 5-minute intervals increased the diversity of traffic situations. Each image was then manually annotated. As the labeling software CVAT [43] lacks native support for parallelogram annotations, we followed the L-shape, triangle-based approach proposed by Fleck et al. [42]: By annotating three outermost vehicle corners, the fourth corner is inferred by reflecting one point across the diagonal, exploiting parallelogram geometry. Since the TUMTraf images provide valuable viewpoint diversity, we selected 282 images from this dataset. Due to their lack of quality, we discarded the original object annotations and manually created new parallelogram labels according to the previously described procedure. Even with the efficient triangle-based annotation approach, manual labeling is resource-intensive, so the size of the real-world dataset is small. Further, the limited number of available roadside cameras restricts the diversity of viewpoints. To enrich the dataset variety, we employ the simulation engine CARLA [44]. We manually placed 818 unique camera positions spread across seven CARLA maps and utilized CARLA-provided ground truth 3D bounding boxes for all spawned vehicles. We obtained the 2.5D ground plane labels by selecting the four bottom vertices, discarding the top vertices. To mitigate bounding box lag, CARLA’s synchronous mode was utilized. Table I summarizes the dataset composition used for training the proposed 2.5D object detector.
@@
Given a monocular image from a single roadside camera, the object detector’s task is to predict the ground plane of all vehicles. 2D/3D object detectors typically leverage their rectangular bounding box structure to regress the center, width, height, and, in the case of 3D, length and rotation. In a similar approach, a parallelogram could be parameterized by six values $x,y,w,h,\theta,\alpha$. Here, $x,y$ denotes the center, $w,h$ the side lengths, $\theta$ the rotation, and $\alpha$ the internal slant angle. Yet $\alpha$ alone is ambiguous, as a parallelogram contains two distinct internal angles, each occurring at a pair of opposite vertices. So we additionally would need to specify the vertex for which $\alpha$ is predicted. To address this limitation, we propose a representation based on the geometric property that a parallelogram defined by vertices $(p_{1},p_{2},p_{3},p_{4})$ and center point $p_{0}$ can be reconstructed if $(p_{0},p_{1},p_{2})$ are known, by mirroring $p_{1}$ and $p_{2}$ across $p_{0}$ to create $p_{3}$ and $p_{4}$. Thus, predicting only $p_{0},p_{1},$ and $p_{2}$ suffices to reconstruct the full parallelogram. Moreover, flipping the points $p_{1}$ and $p_{2}$ does not change the resulting parallelogram. Consequently, our network regresses a triangle composed of the center and the two front corners, rather than the entire parallelogram. Accordingly, the network outputs six values per vehicle detection, specifically the $(x,y)$ coordinates of the points $p_{0}$, $p_{1}$, and $p_{2}$.
@@ III-B2 Loss Function for Regression
  • Pretrain on synthetic data and then subsequently finetune on real-world data.
  • Pretrain on synthetic data and then subsequently finetune on a 70 % real-world / 30 % synthetic mix.
@@ III-D Inference and Deployment
While in training only three vertices are regressed, in inference we obtain the full parallelogram by mirroring the two outer vertices across the predicted center point. The proposed regression head and loss lead to many predictions of overlapping parallelograms because, during training, the one-to-many label assignment maps multiple predictions to the same ground truth. Typically, this is resolved by applying non-maximum suppression (NMS), discarding low-confidence detections that have a high IoU with high-confidence detections. As detailed earlier, calculating IoU for triangles or parallelograms is computationally intensive, so we replace it with an inexpensive approximation during NMS. For each predicted parallelogram, we enclose it in an axis-aligned rectangle derived from its vertices and calculate the IoU for these rectangles, significantly reducing computational complexity. In testing, the average discrepancy between the exact IoU and the approximated IoU was 0.05, which, for reducing total inference time by a factor of three, is an acceptable approximation error.
@@ IV-A Evaluation Metrics
IoU-based metrics are unsuitable for orientation evaluations, as large orientation errors can still lead to high IoU values (cf. Fig. 6). Inspired by 3D object detection metrics, we utilize the mean absolute orientation error (mAOE). For each predicted parallelogram, we take the midpoint $x_{1},y_{1}$ of its front edge and the midpoint $x_{2},y_{2}$ of its rear edge. The same procedure for the ground-truth box yields $\hat{x}_{1},\hat{y}_{1}$ and $\hat{x}_{2},\hat{y}_{2}$. Using these four points, the absolute orientation angle error is calculated by
@@
Figure 6: Visualization of the need of mAOE metric in addition to IoU-based metrics. Predictions in red, annotations in green. IoU based metrics rate the left vehicle prediction as worse, while mAOE metric rates the right vehicle’s prediction as worse.
@@ V Conclusion
In this work, we presented a 2.5D object detection framework tailored to the unique geometry and operating constraints of roadside infrastructure cameras. By reformulating detections as direct predictions of ground-plane parallelograms and implementing this via a triangle-based regression head, we eliminate the need for calibration, road-layout priors, and trajectory tracking that limited earlier approximative projection-based approaches. To compensate for the scarcity and viewpoint bias of existing roadside datasets, we introduced a hybrid training strategy that combines synthetically generated and real-world images. Evaluation on three distinct test datasets shows that our TAF-v2.5 model yields strong generalisation capabilities to unseen viewpoints and weather conditions, while maintaining real-time performance. Compared to the previous TAF-v1 2D-projection baseline, our approach achieves higher localization and orientation accuracy, confirming the benefits of direct ground-plane prediction. Model weights and code are publicly available to encourage future research in intelligent roadside infrastructure perception.
@@ Acknowledgment
This work was supported by funding from the Topic Engineering Secure Systems of the Helmholtz Association (HGF) and by KASTEL Security Research Labs (46.23.03). It was also supported under the ’AI for Mobility’ service contract, funded by the German Ministry of Transport, Baden-Württemberg.

### v2

@@ Abstract
On-board sensors of autonomous vehicles can be obstructed, occluded, or limited by restricted fields of view, complicating downstream driving decisions. Intelligent roadside infrastructure perception systems, installed at elevated vantage points, can provide wide, unobstructed intersection coverage, supplying a complementary information stream to autonomous vehicles via vehicle-to-everything (V2X) communication. However, conventional 3D object-detection algorithms struggle to generalize to the domain shift introduced by top-down perspectives and steep camera angles. We introduce a 2.5D object detection framework, tailored specifically for roadside infrastructure cameras. Unlike conventional 2D or 3D object detection, we employ a prediction approach to detect ground planes of vehicles as parallelograms in the image frame. The parallelogram preserves the planar position, size, and orientation of objects while omitting their height, which is unnecessary for most downstream applications. For training, a mix of real-world and synthetically generated scenes is leveraged. We evaluate generalizability on a held-out camera viewpoint and on adverse-weather scenarios absent from the training set. Our results show high detection accuracy, strong cross-viewpoint generalization, and robustness to diverse lighting and weather conditions. Model weights and inference code are provided at:
@@ I Introduction
Autonomous vehicles primarily rely on on-board sensors such as cameras and Lidar for environmental perception. However, these sensors are inherently limited by occlusions and narrow fields of view, posing challenges in dynamic urban environments where relevant objects may not be visible. This is particularly critical at traffic intersections, where autonomous driving systems must explicitly handle non-observable regions to ensure safe driving maneuvers [1, 2]. Unlike vehicle-mounted sensor systems, roadside infrastructure-based methods leverage elevated vantage points to capture wider fields of view with reduced occlusions and can transmit information to vehicles via V2X communication [3, 4]. However, these elevated viewpoints also introduce new challenges. Although 3D object detection models perform well with on-board vehicle sensors, they often fail to generalize to elevated viewing angles. Furthermore, to reduce costs, intelligent infrastructure typically employs monocular cameras rather than Lidar, limiting the range of applicable 3D object detection methods. While autonomous vehicle perception has abundant datasets [5, 6, 7], infrastructure-based perception datasets are limited, necessitating efficient data collection and training strategies. Current state-of-the-art methods exhibit several limitations, such as assuming uniform vehicle dimensions, enforcing alignment with lane directions and past trajectories, or relying on planar road surfaces [8, 9, 10, 11, 12, 13]. To address these challenges, we present a real-time–capable 2.5D object detection framework. It can be trained efficiently with minimal real-world data, generalizes across diverse camera angles and weather conditions, and outperforms previous approaches.
@@ II-A Camera-based Object Detection
Camera-based object detection is typically categorized into 2D and 3D detection, both aiming to distinguish relevant objects from the background. Standard 2D object detectors detect objects by predicting their $(x,y)$ position, width, and height in pixel coordinates to generate axis-aligned boxes, which are then classified. Most baseline architectures are trained and evaluated on the MS COCO dataset [15] with large architectures, oftentimes based on Detection Transformer (DETR) [16], achieving benchmark performance [17, 18, 19]. A separate line of research prioritizes inference speed to enable real-time object detection without significantly sacrificing accuracy. Here, YOLO-based models [20, 21, 22, 23, 24, 25, 26, 27] and RT-DETR methods [28, 29, 30] are among the most commonly employed approaches.
3D camera-based object detection detects objects in real-world 3D coordinates. These methods typically predict a rectangular cuboid, defined by the object’s $(x,y,z)$ position, width, height, and length. The object’s orientation, expressed in Euler angles, is also predicted. Because most 3D object detection datasets in the traffic domain require additional Lidar data for ground truth annotations, they are typically constructed using data collected from vehicles such as KITTI [5], nuScenes [6], or Waymo [7]. In this domain, the majority of detected objects (vehicles and pedestrians) are situated on a mostly planar road surface, allowing for the common simplification of assuming zero pitch and roll angles. Monocular 3D object detection is an inherently ill-posed problem, as 2D images introduce scale ambiguities that must be approximated by inferring depth and learning geometric priors of frequently encountered objects [31].
One monocular approach, M3D-RPN [32], combines 2D and 3D region proposal networks using 3D anchors to create strong priors correlating 2D scale with 3D depth. For orientation, only the yaw angle is predicted, and it is further refined by reprojecting the 3D cuboid into the image plane and rotating objects to maximize overlap with 2D detections. Another approach, FCOS3D [14], extends the anchor-free FCOS [33] 2D object detection framework to monocular 3D object detection by projecting the 3D object centers onto the image plane and decoupling 2D attributes from depth and orientation estimates.
@@ II-B Object Detectors for Roadside Infrastructure
[8], [9], [10], and [11] rely on 2D object detections projected from image frame into world coordinates. They assume fixed object sizes and estimate orientation from past trajectories. These approaches tackle the ill-posed nature of monocular 3D detection by relying on geometric approximations and precise camera calibration. For example, Rezaei et al. [11] employ approximately 70 images per camera, along with top-view satellite imagery, to perform inverse perspective mapping [34] and project 2D detections into 3D. They assign a fixed width and length to each object class, which is used for depth estimation of 2D object detections. To estimate the object’s heading (yaw angle), they track past detections and approximate the orientation based on the object’s motion trajectory. Since trajectory-based orientation estimation fails for static objects, they also assume parallel heading relative to known road lanes. They evaluate their method on 2D datasets [35, 36, 37] and the non-public Leeds [11] dataset.
@@
Zhu et al. [13], similarly to Rezaei et al. [11], annotate keypoints in satellite and roadside camera images and compute a homography mapping images of roadside cameras into a bird’s-eye view (BEV) assuming planar roads. 2D object detection is subsequently applied to these BEV images, estimating rotated bounding boxes, which correspond to the yaw-angle orientation. However, transforming roadside images into BEV space creates large distortions.
@@ III-A Dataset Creation
For training the proposed 2.5D object detector, datasets containing annotated parallelograms are required. However, to the best of our knowledge, no publicly available dataset is dedicated to ground-plane estimation from roadside cameras. Popular datasets like MIO-TCD [35], UA-DETRAC [36], and GRAM-RTM [37] provide only 2D bounding box information and therefore lack the geometric information needed to recover ground planes. Conversely, 3D datasets are suitable in principle, as the required ground plane information can be calculated by projecting the four bottom vertices of each 3D bounding box into the image plane. Since the Rope3D [39] and Leeds [11] datasets are inaccessible, we examine the TUMTraf [4, 41] dataset for potential use. While camera viewpoints are suitable, we observe that a significant portion of annotations are inaccurate (cf. Figure 4), limiting the dataset’s direct applicability to training our model.
@@
To obtain reliable data, we recorded a new dataset with the intelligent roadside infrastructure of the Test Area Autonomous Driving Baden-Württemberg (TAF-BW) [8, 42, 43]. We captured 3,633 images using seven roadside cameras. Sampling images at 5-minute intervals increased the diversity of traffic situations. Each image was then manually annotated. As the labeling software CVAT [44] lacks native support for parallelogram annotations, we followed the L-shape, triangle-based approach proposed by Fleck et al. [43]: By annotating three outermost vehicle corners, the fourth corner is inferred by reflecting one point across the diagonal, exploiting parallelogram geometry. Since the TUMTraf images provide valuable viewpoint diversity, we selected 282 images from this dataset. Due to their poor annotation quality, we discarded the original object annotations and manually created new parallelogram labels according to the previously described procedure. Even with the efficient triangle-based annotation approach, manual labeling is resource-intensive, so the size of the real-world dataset is small. Furthermore, the limited number of available roadside cameras restricts the diversity of viewpoints. To enrich the dataset variety, we employ the simulation engine CARLA [45]. We manually placed 818 unique camera positions spread across seven CARLA maps and utilized CARLA-provided ground truth 3D bounding boxes for all spawned vehicles. We obtained the 2.5D ground plane labels by selecting the four bottom vertices, discarding the top vertices. To mitigate bounding box lag, CARLA’s synchronous mode was utilized. Table I summarizes the dataset composition used for training the proposed 2.5D object detector.
@@
Given a monocular image from a single roadside camera, the object detector’s task is to predict the ground plane of all vehicles. 2D/3D object detectors typically leverage their rectangular bounding box structure to regress the center, width, height, and, in the case of 3D, length and rotation. In a similar approach, a parallelogram could be parameterized by six values $(x,y),w,h,\theta,\alpha$. Here, $(x,y)$ denotes the center, $w$ and $h$ the side lengths, $\theta$ the rotation, and $\alpha$ the internal slant angle. Yet $\alpha$ alone is ambiguous, as a parallelogram contains two distinct internal angles, each occurring at a pair of opposite vertices. Therefore, we would also need to specify the vertex for which $\alpha$ is predicted. To address this limitation, we propose a representation based on the geometric property that a parallelogram defined by vertices $(p_{1},p_{2},p_{3},p_{4})$ and center point $p_{0}$ can be reconstructed if $(p_{0},p_{1},p_{2})$ are known, by mirroring $p_{1}$ and $p_{2}$ across $p_{0}$ to create $p_{3}$ and $p_{4}$. Thus, predicting only $p_{0},p_{1},$ and $p_{2}$ suffices to reconstruct the full parallelogram. Moreover, flipping the points $p_{1}$ and $p_{2}$ does not change the resulting parallelogram. Consequently, our network regresses a triangle composed of the center and the two front corners, rather than the entire parallelogram. Accordingly, the network outputs six values per vehicle detection, specifically the $(x,y)$ coordinates of the points $p_{0}$, $p_{1}$, and $p_{2}$.
@@ III-B2 Loss Function for Regression
  • Pretrain on synthetic data and subsequently finetune on real-world data.
  • Pretrain on synthetic data and subsequently finetune on a mix of 70 % real-world and 30 % synthetic data.
@@ III-D Inference and Deployment
While in training only three vertices are regressed, in inference we obtain the full parallelogram by reflecting the two predicted non-center vertices across the predicted center point. The proposed regression head and loss lead to many predictions of overlapping parallelograms because, during training, the one-to-many label assignment maps multiple predictions to the same ground truth. Typically, this is resolved by applying non-maximum suppression (NMS), discarding low-confidence detections that have a high IoU with high-confidence detections. As detailed earlier, calculating IoU for triangles or parallelograms is computationally intensive, so we replace it with an inexpensive approximation during NMS. For each predicted parallelogram, we enclose it in an axis-aligned rectangle derived from its vertices and calculate the IoU for these rectangles, significantly reducing computational complexity. In testing, the average discrepancy between the exact IoU and the approximated IoU was 0.05, which, for reducing total inference time by a factor of three, is an acceptable approximation error.
@@ IV-A Evaluation Metrics
IoU-based metrics are inadequate for orientation evaluations, as large orientation errors can still lead to high IoU values (cf. Fig. 6). Inspired by 3D object detection metrics, we utilize the mean absolute orientation error (mAOE). For each predicted parallelogram, we take the midpoint $x_{1},y_{1}$ of its front edge and the midpoint $x_{2},y_{2}$ of its rear edge. The same procedure for the ground-truth box yields $\hat{x}_{1},\hat{y}_{1}$ and $\hat{x}_{2},\hat{y}_{2}$. Using these four points, the absolute orientation angle error is calculated by
@@
Figure 6: Predictions in red, annotations in green. The right vehicle prediction has only a small IoU error to its label. mAOE can be used to still extract the high orientation error. In contrast, the left vehicle prediction has a higher IoU error with small mAOE error.
@@ V Conclusion
In this work, we presented a 2.5D object detection framework adapted to the unique geometry and operating constraints of roadside infrastructure cameras. By reformulating detections as direct predictions of ground-plane parallelograms and implementing this via a triangle-based regression head, we eliminate the need for calibration, road-layout priors, and trajectory tracking that limited earlier approximative projection-based approaches. To compensate for the scarcity and viewpoint bias of existing roadside datasets, we introduced a hybrid training strategy that combines synthetically generated and real-world images. Evaluation on three distinct test datasets shows that our TAF-v2.5 model yields strong generalisation capabilities to unseen viewpoints and weather conditions, while maintaining real-time performance. Compared to the previous TAF-v1 2D-projection baseline, our approach achieves higher localization and orientation accuracy, confirming the benefits of direct ground-plane prediction. Model weights and code are publicly available to encourage future research in intelligent roadside infrastructure perception.
@@ Acknowledgment
This work was supported by funding from the Topic Engineering Secure Systems of the Helmholtz Association (HGF) and by KASTEL Security Research Labs (46.23.03). It was also supported under the ’AI in Mobility’ service contract, on behalf of the Ministry of Transport Baden-Württemberg.